⋆ 65 comparisons Bose and Nelson (1962). ⋆ 63 by Batcher and by Floyd and Knuth (1964). ⋆ 62 by Shapiro (1969) ⋆ 60 by Green (1969)

See the Figure bellow.

Figure 1: Batcher’s proposition of sorting network with 63 comprisons (1964)